Alpha company wants to replace its equipment system (class 8, 20%). The existing equipment can be sold today for $40,000. In another five years it will have a resale value of zero. The new equipment costs $250,000, has a five-year life, and has a $47,000 resale value in five years. The tax rate is 45%, and the opportunity cost of capital is 12%. Due to increased worker productivity and morale, the estimated benefits before tax(∆CFBT) are $65,000 per year. Should the equipment be replaced?
Answer #1

Sale value of existing equipment = 40,000

Cost of New Equipment= 250,000

Resale value of this equipment = 47,000

Tax Rate = 45%

Interest rate/Opportunity cost = 12%

Benefit before tax = 65,000

Now, the decision to replace the equipment will depend on the benefits and cost. If the benefits are more than the cost, the equipment should be replaced.

Total cost= cost of equipment+ Opportunity cost of the capital

= 250,000 + 12% * capital for 5 years

Capital at beginning of year % Interest Interest Capital at the end of year
1st Year 250000 0.12 30000 280000
2nd 280000 0.12 33600 313600
3rd 313600 0.12 37632 351232
4th 351232 0.12 42147.84 393379.8
5th 393379.8 0.12 47205.58 440585.4
Total 190585.4

So, Total cost is 250,000 + 190,585.4=440585, depreciation will not be added as the salvage value is given and depreciation is adjusted there.

Total Benefit: Salvage Value of 1st equipment + Salvage value of second equipment + CFBT after tax for 5 years+ opportunity cost of capital received.

= 40,000+47,000+(65000-45%)*5+ opportunity cost

But, since the opportunity cost is given, the same will also be calculated for 40,000 that we received by selling the 1st equipment and on CFBT after tax every year.

Capital at beginning of year % Interest Interest Capital at the end of year
1st Year 40000 0.12 4800 44800
2nd 44800 0.12 5376 50176
3rd 50176 0.12 6021.12 56197.12
4th 56197.12 0.12 6743.654 62940.77
5th 62940.77 0.12 7552.893 70493.67
Total 30493.67

It is assumed that the equipment was sold at the beginning of the year

Capital at beginning of year % Interest Interest Capital at the end of year
2nd 35750 (65000-45%) 0.12 4290 40040
3rd 75790 (40040+65000-45%) 0.12 9094.8 84884.8
4th 120634.8 (84884.8+65000-45%) 0.12 14476.18 135111
5th 170861 (135111+65000-45%) 0.12 20503.32 191364.3
Total 48364.29

Interest will be earned on CFBT from second year as it is assumed that income is earned at the end of year

No interest will be earned on 47,000 as it is sold at the end of 5 years

Total Benefit= 40000+47000+65000-45%*5+30493.67+48364.29=344608

Since the benefit is less than the cost, equipment should not be replaced.